OpenStudy (anonymous):

find least common multiple x^2-8x, x^2

OpenStudy (raden):

it is x^2

OpenStudy (anonymous):

no sorry its not

OpenStudy (anonymous):

@e.cociuba

OpenStudy (raden):

opps.. forget, i told tha'ts the GCD :) it is x^2-8x

OpenStudy (raden):

ahaa, i remember now :) u havo to factor out the first term x^2-8x = x(x-8) so, the LCM of x(x-8) and x^2 is x^2 * (x-8) = x^3-8x^2

OpenStudy (raden):

do u have options ?

OpenStudy (anonymous):

idk what is the final answer!:?

OpenStudy (anonymous):

it said it was wrong idk